Roselyn/Art- continues work on the MARC Holdings project and believes Artís titles will be completed soon.
Irina/Alex- busy processing reserve material for the summer semester, as well as material she has began receiving for the fall semester.
Shirley/Alex- received 18 carts of donated books, that are being searched in IRIS to determine RUL holdings, findings will be forwarded to the appropriate selectors; also searching items on the quarterly lst/paid, lst/rpl and missing book lists.
Eddie/LSM- reviewed items in Special Collections Room to make sure everything has a record online, some records of periodical titles are not reflecting individual volumes and will need to be updated; progress on the weeding project has slowed down due to lack of space to house items that are pulled; received notification from RU Fire inspector that some ranges of the collection don't adhere to the regulation of a 24 inch height between the top of shelves to the ceiling, a plan of action will be discussed in the coming weeks.
Rob/Alex- finished shifting the problem areas in the stacks; gathering information to create new range guides for areas that were shifted; has caught up on the backlog in shelving of government documents; has begun relabelling books in the Reference collection.
Nicole/Chang- planning a readjustment of shelving in bound periodicals, to accommodate items from new binder that have increased in height.
Megan/Physics- received a substantial donation of books; reviewed and removed some ref-com and media items from permanent reserve and now has more storage space; faulty ceiling tiles have been replaced, as per fire inspector.
Marty/LSM- working on monthly IN-TRANSIT reports and noticing a significant number of items for the ANNEX appearing on them, they may be items from the weeding project.
Brian/Alex- continues work on an online directory of the ALEX collection, planning to coordinate the new paper range guides being made with his online directory.
Matt/Douglass- sorted the folio collection based on a systems report to determine what items are not cataloged as folio, results will be forwarded to systems to update the records; sorting shelves were relocated to create more space around photocopiers.
Jeff/Alex- a section on level BB has been cleared out and will be used as an official NBL storage area.
Roger/NBL- announced NBL staffing changes: Paul Young has been promoted and is now responsible for NBL Government Documents, Jazmine Faherty is now full time staff at ANNEX and Penny Weniger will be returning to manage Douglass.
Andy distributed copies of RULís current policy as well as a document distributed to his staff about enforcing the policy; members were asked to review the policy and think of revisions that would clearly state what is and is not permitted in the libraries.
the policy will need to be realistic to enforce, without being too lenient or to authoritative.
enforcement of the policy is a concern and will require consistent staff cooperation from all units.
patron cooperation is also very important and some type of awareness campaign will be needed.
revisited an ongoing discussion about having designated areas in libraries for eating and drinking; this is something to keep in mind for future consideration if units are renovated.
more ideas will be discussed in future meetings until a final proposal is drafted.
training for CM staff about planning and executing large shifting projects, as well other relevant training sessions led by guest speakers.
incorporate the use of PDAís with scanners into the workflow, to reduce manpower needed for large projects, such as inventory and weeding.
identify and incorporate systems generated reports into the collection management workflow, for correction of problems that may arise.
work with CM staff to incorporate IN-PROCESS statistics into space management procedures, to aid with shifts.
complete weeding and shifting projects across NBL as needed.
site visits of all NBL units to access status of collections.
in searching the missing list for the last quarter, ALEX CM staff noticed a large number of items that were ITEMCAT 2 - RECON. It turns out these were items that had been requested via RDS, not found and charged to missing to have the hold move onto another location.
about 173 + items need to be withdrawn. (It had been determined to accelerate the withdrawal of RECON items since they probably have not been on the shelf for some time.) To handle this volume (and future items), a report will be run a few times per year to identify the items. A final search will be done and the items will be withdrawn by report shortly afterwards. A list will be forwarded to the selectors for their information. Exact timing details are yet to be determined.
Wednesday August 11, 2004